OpenStudy (anonymous):

You would like to have $1,000,000 in 33 years by making regular deposits at the end of each month in an annuity that pays 7% compounded monthly. Determine the deposit at the end of each month. Do not round until the final answer. Then round to the nearest dollar as needed.

OpenStudy (wolf1728):

Okay, I have attached the formula for solving the monthly amount of a monthly annuity.

OpenStudy (wolf1728):

To use the formula the time is the number of months +1 and the rate = the (annual rate / 12) Monthly Amount = 1,000,000 /({ [(1.0058333333)^397 -1]/.0058333333 } -1) Monthly Amount = 1,000,000 / {[(10.065411803 -1) / .0058333333]-1} Monthly Amount = 1,000,000 / ((9.065411803 / .0058333333)-1) Monthly Amount = 1,000,000 / (1,554.0705948003 -1) Monthly Amount = 1,000,000 / 1,553.0705948003 Monthly Amount = 643.89 Here's a monthly annuity calculator to check the calculations http://1728.org/annuitym.htm Then again it should be interest compounded monthly.

OpenStudy (wolf1728):

7 per cent interest compounded annually = 7.229008086 compounded monthly monthly rate = 7.229008086 / 1200 = 0.0060241734 Monthly Amount=1,000,000 /({ [(1.0060241734)^397 -1] / .0060241734}-1) Monthly Amount=1,000,000 /({ [10.8527873492 -1) / .0060241734}-1) Monthly Amount=1,000,000 / ({[9.8527873492] / .0060241734}-1) Monthly Amount=1,000,000 / ({1,635.5417891822}-1) Monthly Amount=1,000,000 / ({1,634.5417891822}) Monthly Amount=611.79 And you can check this amount with the annuity calculator http://1728.org/annuitym.htm Just remember to put 7.229008086 in the yearly rate input box.